OpenStudy (anonymous):

How do you differentiate Y= x squared(1+3x)power 5? Do you use the product rule or chain rule?

OpenStudy (phi):

Use both the product rule and the chain rule let \(u= x^{2}\) and \(v= (1+3x)^{5}\) \[ \frac{d\ u\cdot v}{dx}= v\frac{du}{dx} +u \frac{dv}{dx} \] You can use the chain rule to find \( \frac{dv}{dx}\) let \(w= (1+3x)\) and \(v= w^{5}\) \[ \frac{dv}{dx}= \frac{dv}{dw}\cdot \frac{dw}{dx} \] This gives us \[ \frac{dv}{dx}= 5w^{4} \cdot 3= 15w^{4} \] replacing w with (1+3x): \[ \frac{dv}{dx}= 15(1+3x)^{4} \]